# User-Creator Feature Dynamics in Recommender Systems with Dual Influence

### v1

@@ B.3 Top-$k$ Recommendation
For the experiments on MovieLens 20M dataset, we have $n=2000$ creators and $m=2000$ users, with feature dimension $d=16$. The results for top-$k$ recommendation are in Table 2 and Figure 10. Similar to the experiments with synthetic data, we see that a smaller $k$ improves Creator Diversity (CD) and Recommendation Relevance (RR), reduces Tendency to Polarization (TP), yet worsens Recommendation Diversity (RD).
Table 2: Diversity improvement by top-$k$ recommendation on MovieLens 20M dataset
$\beta$ $k$ CD RD RR TP
$1000$ $0.30\pm 0.04$ $0.03\pm 0.01$ $0.88\pm 0.01$ $1.00\pm 10^{-3}$
$500$ $1.10\pm 0.06$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.43\pm 0.03$
$100$ $1.36\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.28\pm 0.01$
$10$ $1.40\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.20\pm 10^{-3}$
$1$ $2000$ $1.00\pm 10^{-3}$ $\mathbf{0.42\pm 10^{-3}}$ $0.92\pm 0.01$ $1.00\pm 10^{-3}$
$1000$ $0.61\pm 0.16$ $0.03\pm 0.01$ $0.97\pm 0.01$ $0.90\pm 0.06$
$500$ $1.14\pm 0.04$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.41\pm 0.04$
$100$ $1.35\pm 0.01$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.27\pm 10^{-3}$
$10$ $1.40\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.20\pm 10^{-3}$
$3$ $2000$ $0.92\pm 0.07$ $\mathbf{0.02\pm 0.01}$ $0.99\pm 10^{-3}$ $0.91\pm 0.05$
$1000$ $0.65\pm 0.18$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.69\pm 0.14$
$500$ $1.07\pm 0.07$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.48\pm 0.11$
$100$ $1.36\pm 0.01$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.27\pm 0.01$
$10$ $1.40\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.20\pm 10^{-3}$
B.4 Truncation
@@
We also experimented with truncation on synthetic data and MovieLens data. The effect of a large truncation threshold $\tau$ is similar to the effect of a small $k$ in top-$k$ recommendation.
@@
Table 3 shows that the effect of truncation with different thresholds on synthetic data. We see that truncating at $\tau=0$, which corresponds to $90^{\circ}$ angle between $\bm{u}_{j}$ and $\bm{v}_{i}$, is not good for diversity, resulting in the lowest creator diversity measure (CD) and highest tendency to polarization (TP). Truncating at a large threshold like $0.707$ is good for diversity, instead. Figure 11 shows how the diversity measures change over time, under different truncation thresholds.
Results for truncation with different thresholds on the MovieLens 20M dataset are in Table 4 and Figure 12. Similar to synthetic data, we see that a large (but not too large) threshold like 0.707 is good for improving CD and TP.
Table 3: Inner product truncation with different thresholds on synthetic data
@@
Table 4: Inner product truncation with different thresholds on MovieLens 20M dataset
@@
Appendix C Missing Proofs from Section 3
C.1 Proof of Proposition 3.2
We prove a stronger proposition than Proposition 3.2. We say that dynamics $(\bm{U}^{t},\bm{V}^{t})$ reach $(R,\bm{c})$-bi-polarization (or $(R,\bm{c})$-consensus) if the dynamics reaches $R$-bi-polarization (or $R$-consensus) with the vector $\bm{c}$. We prove that $(R,c)$-bi-polarization states are absorbing, which implies that $R$-bi-polarization states are absorbing.
Proof.
Since $\tilde{\bm{v}}_{i}^{t}$ is close to $+\bm{c}$ or $-\bm{c}$, $\langle\tilde{\bm{v}}_{i}^{t},\bm{u}_{j}^{t}\rangle>0$, and $\bm{u}_{j}^{t}$ is close to $+\bm{c}$, it must be that $\tilde{\bm{v}}_{i}^{t}$ is close to $+\bm{c}$, so $\|\tilde{\bm{v}}_{i}^{t}-\bm{c}\|_{2}\leq R$. Then, since $\bm{u}_{j}^{t+1}$ is the normalization of a vector in the convex cone formed by $\bm{u}_{j}^{t}$ and $\tilde{\bm{v}}_{i}^{t}$, by Lemma E.2, we have
We note that every $\tilde{\bm{u}}_{j}^{t}$ satisfies $\|\tilde{\bm{u}}_{j}^{t}-\bm{c}\|_{2}\leq R$ (by the same reasoning as above). Then, since $\bm{v}_{i}^{t+1}$ is the normalization of a vector in the convex cone formed by $\bm{v}_{i}^{t}$ and $\{\tilde{\bm{u}}_{j}^{t}\}_{j\in J}$, by Lemma E.2, we have
C.2 Proof of Theorem 3.3
The proof of Theorem 3.3 uses the following lemma:
Lemma C.1.
Proof of Theorem 3.3.
For any state $(\bm{U}^{t},\bm{V}^{t})$ in the state space, by Lemma C.1 there exists a path $(\bm{U}^{t},\bm{V}^{t})\to\cdots\to(\bm{U}^{t+T},\bm{V}^{t+T})$ of length $T$ that leads to $R$-bi-polarization. Because every creator can be recommended to a user with probability at least $p_{0}$, each transition $(\bm{U}^{t^{\prime}},\bm{V}^{t^{\prime}})\to(\bm{U}^{t^{\prime}+1},\bm{V}^{t^{\prime}+1})$ happens with probability at least $p_{0}^{m}$. So, the path of length $T$ has probability at least $p_{0}^{mT}>0$, and the probability that the dynamics does not reach $R$-bi-polarization after $KT$ steps is at most $(1-p_{0}^{mT})^{K}$, which $\to 0$ as $K\to\infty$. Therefore, with probability $1$ the dynamics will reach $R$-bi-polarization eventually. ∎
Appendix D Missing Proofs from Section 4
D.1 Proof of Proposition 4.1
Before proving Proposition 4.1, we formally define what “clusters” mean:
Definition D.1 (clusters).
Proof of Proposition 4.1.
$\|\bm{u}_{j}^{t}-\bm{v}_{i}^{t}\|\leq 2R.$
$\|\bm{u}_{j}^{t}-\bm{v}_{i^{\prime}}^{t}\|>2R$
Since $B(\bm{c}_{\ell},R)$ contains $k$ creators, these $k$ creators are the $k$-most relevant ones to user $j$, so user $j$ will only be recommended these creators. Then, by applying Proposition 3.2 to each of the $\lfloor n/k\rfloor$ balls separately, we see that each ball is a $R$-consensus and hence absorbing. So, the new state $(\bm{U}^{t+1},\bm{V}^{t+1})$ still forms $\lfloor n/k\rfloor$ clusters with these $\lfloor n/k\rfloor$ balls. ∎
Appendix E Useful Lemmas
@@
Lemma E.2 (Convex Cone Property).
@@
This proves the first inequality. To prove the second inequality, we use Claim E.1 and the first inequality:
@@
Because $\langle\bm{x}^{t},\bm{z}^{t}\rangle\geq 0$, the vector $\tilde{\bm{x}}^{t+1}=\bm{x}^{t}+\eta\bm{z}^{t}$ has length $\geq 1$ and hence is outside (or on the surface) of the $d$-dimensional unit ball. Since $\bm{x}^{t}=\mathcal{P}(\tilde{\bm{x}}^{t+1})$ is the projection of $\tilde{\bm{x}}^{t+1}$ onto the unit ball, and $\bm{z}^{t}$ is another vector inside the unit ball, by the “Pythagorean property” (Proposition 2.2 in (Bansal & Gupta, 2019)), we must have $\langle\bm{x}^{t}-\bm{x}^{t+1},\tilde{\bm{x}}^{t+1}-\bm{x}^{t+1}\rangle\leq 0$. This implies
@@
For $\langle\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle$, using Lemma E.3 with $\bm{x}^{t}=\bm{u}_{j}^{t},\bm{z}^{t}=\bm{v}_{i}^{t}$, and $\eta=\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})$, we get
@@
For $\langle\bm{u}_{j}^{t+1},\bm{v}_{i}^{t+1}-\bm{v}_{i}^{t}\rangle$, by Cauchy-Schwarz inequality and Lemma E.4,
@@
We prove the first item by induction. Suppose $\langle\bm{u}_{j}^{t},\bm{y}\rangle>\langle\bm{v}_{i}^{t},\bm{y}\rangle>0$ holds. Consider $t+1$. First, by Lemma E.2, $\langle\bm{v}_{i}^{t+1},\bm{y}\rangle>0$ holds. Then, we prove $\langle\bm{u}_{j}^{t+1},\bm{y}\rangle>\langle\bm{v}_{i}^{t+1},\bm{y}\rangle$. Let $f=f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})$.
@@
We then prove the second item. Using Lemma E.3 for $\bm{v}_{i}^{t+1}=\mathcal{P}(\bm{v}_{i}^{t}+\eta_{c}\bm{u}_{j}^{t})$, we get
@@
Using Lemma E.3 for $\bm{u}_{j}^{t+1}=\mathcal{P}(\bm{u}_{j}^{t}+\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})\bm{v}_{i}^{t})$ and using the fact $\langle\bm{v}_{i}^{t},\bm{y}\rangle-\langle\bm{u}_{j}^{t},\bm{y}\rangle<0$ proved in item 1,
@@
Appendix F Proof of Lemma C.1
Lemma C.1 is proved by induction on the number $n$ of creators. We first show that any system with 1 creator and multiple users must converge to $R$-bi-polarization in finite steps for any $R>0$. Using the result for $1$ creator, we then construct a finite length path that leads to $R$-bi-polarization for any system with $n\geq 2$ creators.
@@ F.1 Base Case: Convergence Results for $n=1$ Creator
We first note that, by Lemma E.5, all user vectors satisfy $\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle>0$ in all time steps $t>0$. Hence, the creator update is always $\bm{v}_{i}^{t+1}=\mathcal{P}(\bm{v}_{i}^{t}+\frac{\eta_{c}}{|J|}\sum_{j\in J}g(\bm{u}_{j}^{t},\bm{v}_{i}^{t})\bm{u}_{j}^{t})=\mathcal{P}(\bm{v}_{i}^{t}+\eta_{c}\frac{1}{|J|}\sum_{j\in J}\bm{u}_{j}^{t})$.
@@
Using Lemma E.3 with $\bm{x}^{t}=\bm{u}_{j}^{t}$, $\bm{z}^{t}=\bm{v}_{i}^{t}$, and $\eta=\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})$, we get
@@
Using Lemma E.4 with $\bm{x}^{t}=\bm{u}_{j}^{t}$, $\bm{z}^{t}=\bm{v}_{i}^{t}$, and $\eta=\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})$, we get
@@
Using Lemma E.4 with $\bm{x}^{t}=\bm{v}_{i}^{t}$, $\bm{z}^{t}=\frac{1}{|J|}\sum_{j\in J}\bm{u}_{j}^{t}$, and $\eta=\eta_{c}$, we get
@@
Let $J^{+}=\{j\in J:\langle\bm{u}_{j}^{0},\bm{v}_{i}^{0}\rangle>0\}$ be the set of users with positive inner products with creator $i$ initially; let $J^{-}=\{j\in J:\langle\bm{u}_{j}^{0},\bm{v}_{i}^{0}\rangle<0\}$. Let $\tilde{\bm{u}}_{j}^{t}=-\bm{u}_{j}^{t}$ for $j\in J^{-}$ and $\tilde{\bm{u}}_{j}^{t}=\bm{u}_{j}^{t}$ for $j\in J^{+}$. Then, the system consisting of $\{\tilde{\bm{u}}_{j}^{t}\}_{j\in J}$ and $\bm{v}_{i}^{t}$ satisfies the initial condition $\langle\tilde{\bm{u}}_{j}^{0},\bm{v}_{i}^{0}\rangle>0$ in Lemma F.1. So, by Lemma F.1, it reaches $R$-consensus after at most $\frac{8}{3\eta_{u}L_{f}}\ln\frac{2|J|}{R^{2}}$ steps. Then by the reflection lemma (Lemma E.7), the original system, consisting of $\{\bm{u}_{j}^{t}\}_{j\in J}$ and $\bm{v}_{i}^{t}$, must reach $R$-bi-polarization. ∎
F.2 Inductive Step: Proof of Lemma C.1
@@
Step (1): Consider the subsystem of the first $n-1$ creators and all users $J$. By induction, there exists a path of length $T_{1}=L_{n-1,R_{1}}<+\infty$ that leads the subsystem to $(R_{1},\bm{c}^{T_{1}})$-consensus with some $\bm{c}^{T_{1}}\in\mathbb{S}^{d-1}$. So, after these $T_{1}$ steps, all creators $i\in\{1,\ldots,n-1\}$ and all users $j\in J$ satisfy $\|\bm{v}_{i}^{T_{1}}-\bm{c}^{T_{1}}\|\leq R_{1}$ and $\|\bm{u}_{j}^{T_{1}}-\bm{c}^{T_{1}}\|\leq R_{1}$. Creator $n$ does not update during these $T_{1}$ steps, so $\bm{v}_{n}^{T_{1}}=\bm{v}_{n}^{0}$, and it still has positive inner products with the first $n-1$ creators and all users by the convex cone property (Lemma E.2). Let’s then consider the distance between creators $n$ and the consensus center $\bm{c}^{T_{1}}$: $\|\bm{v}_{n}^{T_{1}}-\bm{c}^{T_{1}}\|$. If $\|\bm{v}_{n}^{T_{1}}-\bm{c}^{T_{1}}\|\leq R$, then the system has satisfied $(R,\bm{c}^{T_{1}})$-consensus, so our construction is finished. Otherwise, $\|\bm{v}_{n}^{T_{1}}-\bm{c}^{T_{1}}\|>R$. We continue the construction as follows:
@@
Thus, we can apply Lemma E.6 with $\bm{y}=\bm{c}^{T_{1}}$ to derive that, after these $T_{2}$ steps,

### v2

@@ 5.3 Top-$k$ Truncation and Threshold Truncation
We also experimented with threshold truncation on synthetic data (Appendix B.2) and MovieLens data (Appendix C). The effect of a large truncation threshold $\tau$ is similar to the effect of a small $k$ in top-$k$ truncation.
6 Conclusion
Our work defines a dynamics model to capture the dual influence of recommender systems on user preferences and content creation. Although our model is a theoretical abstraction, we believe that it captures the essence of a real-world recommender system, and our effort is an important initial endeavor to study diversity in recommender systems with dual influence. (See Appendix H for some additional discussions on real-world recommender systems.) We specifically point out different concepts of diversity in recommender systems (creation diversity, recommendation diversity, and tendency to polarization) and provide theoretical and empirical evidences to show that, due to dual influence, myopically optimizing recommendation diversity might hurt the long-term creation diversity and result in polarization of the system. We also explore popular design choices in recommender systems and show an interesting and somewhat counter-intuitive result that designs purely targeting efficiency improvement (e.g., top-$k$ truncation) can alleviate polarization. We believe that the insights from our work are valuable to building healthy and sustainable recommender systems, and our results can inspire more sophisticated solutions for improving the long-term diversity of recommender systems to be developed.
Appendix A Additional Discussions on Related Works
Table 2: Comparison between our work and some previous works on performative effects of recommender systems
Works
Adaptive
Users?
Adaptive
Creators?
Creator Reward
Dynamics or
Equilibrium?
Content Adjustment Model
Ours Yes Yes User engagement Dynamics
Conditioned on previous time step;
implicit cost of content adjustment
[15] No Yes Exposure Dynamics
Conditioned on previous time step;
explicit cost of content adjustment
[42] No Yes User engagement Dynamics Freely choose without cost
[35] No Yes User engagement Dynamics Freely choose without cost
[23] No Yes Exposure Equilibrium Freely choose with cost
[20] No Yes Exposure Equilibrium Freely choose without cost
[7] No Yes Exposure Equilibrium Freely choose without cost
[2] No Yes User engagement Equilibrium Freely choose without cost
[43] No Yes
Designed by a welfare-
maximizing platform
Dynamics Freely choose without cost
[14] Yes No11footnotemark: 1 N/A Dynamics N/A
[41] Yes No11footnotemark: 1 N/A Dynamics N/A
[3]
Adaptive and
adversarial
No11footnotemark: 1 N/A Dynamics N/A
11footnotemark: 1
: These works study the design of recommendation algorithms for the platform with a fixed set of content, without explicitly modeling the content creators.
Appendix B Additional Experiments on Synthetic Data
B.1 Number of Fixed Dimensions
In this part, we consider the case where some dimensions of the user feature vectors are fixed features and thus not updated from round to round (e.g., ages, genders). Formally, we fix the first $k\leq d$ dimensions. The remaining $d-k$ dimensions $\bm{u}_{j}^{t}[k+1:d]=(u_{j}^{t}[k+1],\ldots,u_{j}^{t}[d])$ are updated according to the following rule: $\bm{u}_{j}^{t+1}[k+1:d]=\|\bm{u}_{j}^{t}[k+1:d]\|\cdot\mathcal{P}\big{(}\bm{u}_{j}^{t}[k+1:d]+\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})\bm{v}_{i}^{t}[k+1:d]\big{)}$. The multiplication by $\|\bm{u}_{j}^{t}[k+1:d]\|$ ensures unit norm $\|\bm{u}_{j}^{t+1}\|=1$. The effect of the number of fixed dimensions on the dynamics is shown in Figure 7. We see that the diversity of the system improves as the number of fixed dimensions increases, and the degree of polarization is reduced. This is similar to the effect of decreasing user update rate $\eta_{u}$ in Figure 4. The observation that fixed user features encourage diversity might be a reason why the recommender systems in practice are not as polarized as our theoretical prediction.
B.2 Threshold Truncation
@@
Table 3 shows that the effect of different thresholds in threshold truncation on the long-term diversity of the system. We see that truncating at $\tau=0$, which corresponds to $90^{\circ}$ angle between $\bm{u}_{j}$ and $\bm{v}_{i}$, is not good for diversity, resulting in the lowest creator diversity measure (CD) and highest tendency to polarization (TP). Truncating at a large threshold like $0.707$ is good for diversity, instead.
Figure 8 shows how the diversity measures change over time, under different truncation thresholds.
Table 3: Diversity improvement by threshold truncation on synthetic data
@@
Appendix C Additional Experiments on Real-World Dataset
C.1 Details of the Recommendation Algorithm
Algorithm 1 Real-world Recommendation with Dual Influence
  repeat
     Let temporary parameter $\bm{w}^{(0)}\leftarrow\bm{\theta}^{(t)}$
     Compute loss $\mathcal{L}(\bm{\theta}^{(t)},Y^{(t)})$
     for $s=1$ to $m-1$ do
     end for
     $\bm{\theta}^{(t+1)}\leftarrow\bm{w}^{(m)}$
     Deliver recommendations based on $\hat{U}^{(t+1)},\hat{V}^{(t+1)}$
     Update $U^{(t+1)},V^{(t+1)}$, and $Y^{(t+1)}$
     $t\leftarrow t+1$
  until $\|\bm{\theta}^{(t)}-\bm{\theta}^{(t-1)}\|_{2}\leq\delta$
C.2 Top-$k$ and Threshold Truncations
We also try top-$k$ and threshold truncations (Section 4) on the MovieLens 20M dataset. Here, we have $n=2000$ creators and $m=2000$ users, with feature dimension $d=16$. The results for top-$k$ truncation are in Table 4 and Figure 10. Similar to the experiments with synthetic data, we see that a smaller $k$ improves Creator Diversity (CD) and Recommendation Relevance (RR), reduces Tendency to Polarization (TP), yet worsens Recommendation Diversity (RD).
Table 4: Diversity improvement by top-$k$ truncation on MovieLens 20M dataset
$\beta$ $k$ CD RD RR TP
$1000$ $0.30\pm 0.04$ $0.03\pm 0.01$ $0.88\pm 0.01$ $1.00\pm 10^{-3}$
$500$ $1.10\pm 0.06$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.43\pm 0.03$
$100$ $1.36\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.28\pm 0.01$
$10$ $1.40\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.20\pm 10^{-3}$
$1$ $2000$ $1.00\pm 10^{-3}$ $\mathbf{0.42\pm 10^{-3}}$ $0.92\pm 0.01$ $1.00\pm 10^{-3}$
$1000$ $0.61\pm 0.16$ $0.03\pm 0.01$ $0.97\pm 0.01$ $0.90\pm 0.06$
$500$ $1.14\pm 0.04$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.41\pm 0.04$
$100$ $1.35\pm 0.01$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.27\pm 10^{-3}$
$10$ $1.40\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.20\pm 10^{-3}$
$3$ $2000$ $0.92\pm 0.07$ $\mathbf{0.02\pm 0.01}$ $0.99\pm 10^{-3}$ $0.91\pm 0.05$
$1000$ $0.65\pm 0.18$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.69\pm 0.14$
$500$ $1.07\pm 0.07$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.48\pm 0.11$
$100$ $1.36\pm 0.01$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.27\pm 0.01$
$10$ $1.40\pm 10^{-3}$ $0.00\pm 10^{-3}$ $1.00\pm 10^{-3}$ $0.20\pm 10^{-3}$
Results for threshold truncation are in Table 5 and Figure 11. Similar to synthetic data, we see that a large (but not too large) threshold like 0.707 is good for improving CD and TP.
Table 5: Threshold truncation with different thresholds on MovieLens 20M dataset
@@ C.2 Top-$k$ and Threshold Truncations
Appendix D Useful Lemmas
@@
Lemma D.2 (Convex Cone Property).
@@
This proves the first inequality. To prove the second inequality, we use Claim D.1 and the first inequality:
@@
Because $\langle\bm{x}^{t},\bm{z}^{t}\rangle\geq 0$, the vector $\tilde{\bm{x}}^{t+1}=\bm{x}^{t}+\eta\bm{z}^{t}$ has length $\geq 1$ and hence is outside (or on the surface) of the $d$-dimensional unit ball. Since $\bm{x}^{t}=\mathcal{P}(\tilde{\bm{x}}^{t+1})$ is the projection of $\tilde{\bm{x}}^{t+1}$ onto the unit ball, and $\bm{z}^{t}$ is another vector inside the unit ball, by the “Pythagorean property” (Proposition 2.2 in [6]), we must have $\langle\bm{x}^{t}-\bm{x}^{t+1},\tilde{\bm{x}}^{t+1}-\bm{x}^{t+1}\rangle\leq 0$. This implies
@@
For $\langle\bm{u}_{j}^{t+1}-\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle$, using Lemma D.3 with $\bm{x}^{t}=\bm{u}_{j}^{t},\bm{z}^{t}=\bm{v}_{i}^{t}$, and $\eta=\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})$, we get
@@
For $\langle\bm{u}_{j}^{t+1},\bm{v}_{i}^{t+1}-\bm{v}_{i}^{t}\rangle$, by Cauchy-Schwarz inequality and Lemma D.4,
@@
We prove the first item by induction. Suppose $\langle\bm{u}_{j}^{t},\bm{y}\rangle>\langle\bm{v}_{i}^{t},\bm{y}\rangle>0$ holds. Consider $t+1$. First, by Lemma D.2, $\langle\bm{v}_{i}^{t+1},\bm{y}\rangle>0$ holds. Then, we prove $\langle\bm{u}_{j}^{t+1},\bm{y}\rangle>\langle\bm{v}_{i}^{t+1},\bm{y}\rangle$. Let $f=f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})$.
@@
We then prove the second item. Using Lemma D.3 for $\bm{v}_{i}^{t+1}=\mathcal{P}(\bm{v}_{i}^{t}+\eta_{c}\bm{u}_{j}^{t})$, we get
@@
Using Lemma D.3 for $\bm{u}_{j}^{t+1}=\mathcal{P}(\bm{u}_{j}^{t}+\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})\bm{v}_{i}^{t})$ and using the fact $\langle\bm{v}_{i}^{t},\bm{y}\rangle-\langle\bm{u}_{j}^{t},\bm{y}\rangle<0$ proved in item 1,
@@
Appendix E Proof of Proposition 3.2
Proof.
Since $\tilde{\bm{v}}_{i}^{t}$ is close to $+\bm{c}$ or $-\bm{c}$, $\langle\tilde{\bm{v}}_{i}^{t},\bm{u}_{j}^{t}\rangle>0$, and $\bm{u}_{j}^{t}$ is close to $+\bm{c}$, it must be that $\tilde{\bm{v}}_{i}^{t}$ is close to $+\bm{c}$, so $\|\tilde{\bm{v}}_{i}^{t}-\bm{c}\|_{2}\leq R$. Then, since $\bm{u}_{j}^{t+1}$ is the normalization of a vector in the convex cone formed by $\bm{u}_{j}^{t}$ and $\tilde{\bm{v}}_{i}^{t}$, by Lemma D.2, we have
We note that every $\tilde{\bm{u}}_{j}^{t}$ satisfies $\|\tilde{\bm{u}}_{j}^{t}-\bm{c}\|_{2}\leq R$ (by the same reasoning as above). Then, since $\bm{v}_{i}^{t+1}$ is the normalization of a vector in the convex cone formed by $\bm{v}_{i}^{t}$ and $\{\tilde{\bm{u}}_{j}^{t}\}_{j\in J}$, by Lemma D.2, we have
Appendix F Proof of Lemma 3.4
Lemma 3.4 is proved by induction on the number $n$ of creators. We first show that any system with 1 creator and multiple users must converge to $R$-bi-polarization in finite steps for any $R>0$. Using the result for $1$ creator, we then construct a finite length path that leads to $R$-bi-polarization for any system with $n\geq 2$ creators.
@@ F.1 Base Case: Convergence Results for $n=1$ Creator
We first note that, by Lemma D.5, all user vectors satisfy $\langle\bm{u}_{j}^{t},\bm{v}_{i}^{t}\rangle>0$ in all time steps $t>0$. Hence, the creator update is always $\bm{v}_{i}^{t+1}=\mathcal{P}(\bm{v}_{i}^{t}+\frac{\eta_{c}}{|J|}\sum_{j\in J}g(\bm{u}_{j}^{t},\bm{v}_{i}^{t})\bm{u}_{j}^{t})=\mathcal{P}(\bm{v}_{i}^{t}+\eta_{c}\frac{1}{|J|}\sum_{j\in J}\bm{u}_{j}^{t})$.
@@
Using Lemma D.3 with $\bm{x}^{t}=\bm{u}_{j}^{t}$, $\bm{z}^{t}=\bm{v}_{i}^{t}$, and $\eta=\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})$, we get
@@
Using Lemma D.4 with $\bm{x}^{t}=\bm{u}_{j}^{t}$, $\bm{z}^{t}=\bm{v}_{i}^{t}$, and $\eta=\eta_{u}f(\bm{v}_{i}^{t},\bm{u}_{j}^{t})$, we get
@@
Using Lemma D.4 with $\bm{x}^{t}=\bm{v}_{i}^{t}$, $\bm{z}^{t}=\frac{1}{|J|}\sum_{j\in J}\bm{u}_{j}^{t}$, and $\eta=\eta_{c}$, we get
@@
Let $J^{+}=\{j\in J:\langle\bm{u}_{j}^{0},\bm{v}_{i}^{0}\rangle>0\}$ be the set of users with positive inner products with creator $i$ initially; let $J^{-}=\{j\in J:\langle\bm{u}_{j}^{0},\bm{v}_{i}^{0}\rangle<0\}$. Let $\tilde{\bm{u}}_{j}^{t}=-\bm{u}_{j}^{t}$ for $j\in J^{-}$ and $\tilde{\bm{u}}_{j}^{t}=\bm{u}_{j}^{t}$ for $j\in J^{+}$. Then, the system consisting of $\{\tilde{\bm{u}}_{j}^{t}\}_{j\in J}$ and $\bm{v}_{i}^{t}$ satisfies the initial condition $\langle\tilde{\bm{u}}_{j}^{0},\bm{v}_{i}^{0}\rangle>0$ in Lemma F.1. So, by Lemma F.1, it reaches $R$-consensus after at most $\frac{8}{3\eta_{u}L_{f}}\ln\frac{2|J|}{R^{2}}$ steps. Then by the reflection lemma (Lemma D.7), the original system, consisting of $\{\bm{u}_{j}^{t}\}_{j\in J}$ and $\bm{v}_{i}^{t}$, must reach $R$-bi-polarization. ∎
F.2 Inductive Step: Proof of Lemma 3.4
@@
Step (1): Consider the subsystem of the first $n-1$ creators and all users $J$. By induction, there exists a path of length $T_{1}=L_{n-1,R_{1}}<+\infty$ that leads the subsystem to $(R_{1},\bm{c}^{T_{1}})$-consensus with some $\bm{c}^{T_{1}}\in\mathbb{S}^{d-1}$. So, after these $T_{1}$ steps, all creators $i\in\{1,\ldots,n-1\}$ and all users $j\in J$ satisfy $\|\bm{v}_{i}^{T_{1}}-\bm{c}^{T_{1}}\|\leq R_{1}$ and $\|\bm{u}_{j}^{T_{1}}-\bm{c}^{T_{1}}\|\leq R_{1}$. Creator $n$ does not update during these $T_{1}$ steps, so $\bm{v}_{n}^{T_{1}}=\bm{v}_{n}^{0}$, and it still has positive inner products with the first $n-1$ creators and all users by the convex cone property (Lemma D.2). Let’s then consider the distance between creators $n$ and the consensus center $\bm{c}^{T_{1}}$: $\|\bm{v}_{n}^{T_{1}}-\bm{c}^{T_{1}}\|$. If $\|\bm{v}_{n}^{T_{1}}-\bm{c}^{T_{1}}\|\leq R$, then the system has satisfied $(R,\bm{c}^{T_{1}})$-consensus, so our construction is finished. Otherwise, $\|\bm{v}_{n}^{T_{1}}-\bm{c}^{T_{1}}\|>R$. We continue the construction as follows:
@@
Thus, we can apply Lemma D.6 with $\bm{y}=\bm{c}^{T_{1}}$ to derive that, after these $T_{2}$ steps,